OpenStudy (anonymous):

43m+76/7=40/7 (fraction eqation)

OpenStudy (anonymous):

m=-36/43

OpenStudy (aroub):

43m+76/7 =40/7 ok the first step u u multiply 7 by the whole equation to get rid of the 7 in the denominator so it becomes 43m+76=40 , 43m=-76+40 , 43m=-36 m=-36/43